# Employing Sentence Space Embedding for Classification of Data Stream from Fake News Domain

### v1

@@ 4.1 Set-up
Data All of the research was conducted using Fakeddit’s multimodal dataset, which presents a real-life fake news classification task broken down into two, three or six classes [28]. This dataset consists of more than one million posts on 22 different subreddits of the social networking platform Reddit and includes text and image modalities, supplemented by metadata about the posts and their authors. For the purposes of this study, a single binary data stream was prepared, in which consecutive texts were sorted accordingly to their creation timestamp. Of the entire dataset, 682,996 multimodal samples were used (the rest have only one modality). The decision to limit to multimodal samples only is linked to the facilitation of extending the presented research to include the image modality. The data stream was divided into 2731 data chunks, containing 250 samples each. Relevant to the research course is the fact that the resulting data stream is characterized by a dynamic imbalance ratio, which changes in successive batches to the point where, at about 3/4 of the length of the stream, a minority class transitions into a majority class. The exact changes in the prior probability of class membership are shown in Fig. 3.
@@
Experimental protocol & reproducibility All experiments were carried out using the Test-Then-Train protocol to guarantee a robust experimental evaluation, were implemented in Python and can be replicated using the publicly available GitHub repository [^1] https://github.com/w4k2/sentence-space-stream. Implementation of state-of-the-art algorithms were based on stream-learn [18], scikit-multiflow [27], and PyTorch [30] libraries. The classification quality evaluation of the algorithms was based on the standard metrics used in the task of imbalanced data classification, i.e. balanced accuracy score (bac), recall, specificity, precision, $F_{1}$ score, $Gmean$, and $Gmean_{s}$.
@@ 4.2 Experiment scenarios
Experiment 1 – Extraction methods The goal of Experiment 1 was to compare the performance of sss depending on the type of extractor used to obtain a representation for Sentence Space encoding. For this purpose, (i) GloVe, (ii) MiniLM, (iii) pre-trained Word2Vec, and (iv) Word2Vec updated after each data chunk were compared with each other. The representation width depending on the extractor was $380$ for MiniLM and $300$ for GloVe and W2V. Based on the results, the feature extraction method used in subsequent experiments was selected.
Experiment 2 – Comparison with data stream classification algorithms In Experiment 2, the sss based on the extractor chosen in Experiment 1 was compared with state-of-the-art ensemble algorithms for imbalanced data stream classification. Among the methods mentioned in the literature review, (i) Hoeffding Tree with Hellinger split criterion (hf), (ii) Learn++.CDS (cds), (iii) Learn++.NIE (nie), (iv) Kappa Updated Ensemble (kue), and (v) Robust Online Self-Adjusting Ensemble (rose) were selected as references. hf was used as the base classifier for all reference methods, and the maximum size of the classifier pool was set to $10$. The selection of methods, base classifier, and pool size was based on the literature [1, 6, 7]. The entire set of reference algorithms was compared with the SSS depending on the approach used to extract features from the text. In addition to the extractors used in Experiment 1, tf-idf with unigrams and bigrams and 100 features with top term frequency was employed here. A set of reference methods trained using the representation that provided the highest classification quality in terms of bac was selected for the last experiment.
@@ 4.3 Experiment 1 – Extraction methods
As can be seen in Figure 4, across all data chunks, the results of the first three methods GloVe, MiniLM, pre-trained Word2Vec are similar, with the only deviating method being partial-fit Word2Vec trained in each chunk. This discrepancy can be explained by the dictionary’s limitations built on the training data, consisting of article titles, primarily short texts. In contrast, the MiniLM model and Word2Vec pre-trained (word2vec-google-news-300) based on the Google News dataset, as well as GloVe (glove-wiki-gigaword-300), which means they were all built on larger volumes of data compared to the own trained Word2Vec.
Following this experiment, it was decided to use the GloVe vectors for further research, which is a newer method than Word2Vec that processes additional global information and, at the same time – unlike MiniLM, dedicated to sequential processing – was designed to process single words. In addition, the MiniLM model requires more computational complexity when determining the representation vector for a given word, which for Word2Vec and Glove methods is limited only to reading values from the vector array. With equal effectiveness, the exclusion of MiniLM is fully justified.
@@ 4.4 Experiment 2 – Comparison with data stream classification algorithms
The preliminary part of main comparative experiment has begun from analysis of feature space reduction influence. It was conducted by Principal Component Analysis (pca) – projecting the original embedding space down into 100 features. As it can be observed in Figure 5, influence of this simple reduction is mostly cosmetic, always laying in one percent margin of a difference, so it is justified to reduce problem representation for canonical models, since representation gets smaller while change in quality is negligible.
Figure 5: Results of an experiment to decide whether to use PCA for dimensionality reduction of representations for reference methods.
The main comparison of recognition efficiency, taking into account the deep strategy already proposed in this work (sss), is presented in Figure 6, divided into four empirical analyzes according to the extraction strategies used for canonical methods. In the case of each approach, the inglorious laggard turns out to be the nie method, which maintains its model at the level of a random classifier for a very long time, and in two cases (Glove and MiniLM) raises it only slightly in the final phase of the stream, after the prior concept drift shown in Fig. 3.
@@
The weakest method in the main rate (excluding nie) turns out to be Learning++CDS, which for all extractors except MiniLM clearly leans towards randomness. All analyzed reference solutions appear to be very sensitive to concept drift occurring around the chunk 2,000. The ranking of kue, rose and raw Hoeffding Tree methods depends on the extraction strategy used, but the overall distribution of their effectiveness is rather similar, with a slight advantage for rose.
sss as a deep learning strategy shows a noticeable advantage over all canonical solutions throughout the entire data stream, being the only one in the competition to maintain an average recognition efficiency of 80 percent of balanced accuracy score. This observation is supported by the extended analysis of metrics in the form of radar (Figure 7), where an advantage of sss can be observed in each of the simple and aggregate metrics used. The outlier specificity result for the nie method results solely from its complete inability to learn in the analyzed problem environment, which for the dominant majority of processing time induces decisions towards one of the problem classes, without a generalized relation to given bias.
@@
Among the analyzed extraction methods, MiniLM comes minimaly to the fore, as it was the only one that allowed the nie strategy to noticeably rise from the random classifier level in the final part of the stream, and the cds to compete with the rest of the competition.
@@ 4.5 Experiment 3 – Time complexity
The results of the third experiment – showing the time complexity for extracting, training, and testing the algorithm – are shown in Figure 8. As we can see, in the case of preprocessing, all reference methods show uniform computation time (lines overlap), and only the proposed approach deviates from this tendency and performs extraction faster. Additionally, it should be noted that this happens despite the time measurement considering the transition to image representation. The method owes it using the GloVe technique for sss, which in Experiment 1 had the best results for set space encoding. The reference methods, however, use the MiniLM transformer in accordance with the outcome of Experiment 2.
More variability can be observed in the training and testing processes – in both, a single Hoeffding Tree processes the fastest. In turn, training in a single epoch is the slowest for rose, and prediction – for nie. It is also reflected in the accumulated time graph, which indicates that the only method ahead of sss in the entire processing is the Hoeffding Tree. Still, it should be borne in mind that a single classifier is considered here, not an ensemble.
Therefore, the proposed method is ahead of classifiers ensembles used for data streams even though it is based on convolutional networks and requires transformation from text into an image. At the same time, despite the lowest processing time compared to state-of-the-art ensemble algorithms, sss offers the highest generalization ability. All this means that sss can be successfully used in real-life batch-based data stream classification tasks.
5 Conclusion
The presented research work aimed to achieve two main goals. The first was to address the application of deep learning in the task of data stream classification, which is presented in the current literature as one of the main research areas in the need of further investigation. The second goal was to propose – for the first time in the literature – the use of sentence space, which is the equivalent of multi-dimensional encoding for text data, in a data stream classification task.
To realize the above goals, Streaming Sentence Space (sss) was proposed, which encodes the text found in individual data batches into discrete digital signals based on embeddings obtained through the GloVe technique. The resulting images are then classified using the ResNet-18 architecture, which, in order to reduce computational complexity, performs only a single training epoch on each data chunk.
The developed approach was tested on the basis of computer experiments conducted on a real-life dynamically imbalanced data stream formed by chronologically ordering the texts contained in the Fakeddit dataset. The results showed that sss, thanks to the inherent generalization ability of the convolutional neural network, is able to outperform the classification quality of state-of-the-art classifier ensemble methods dedicated for imbalanced data stream classification. In addition, sss exhibits lower time complexity than ensemble reference methods, which further encourages its use and contradicts the popular opinion that deep learning has too high time and computational complexity to be used for data stream analysis.

### v2

@@ 4.1 Set-up
Data All of the research was conducted using Fakeddit’s multimodal dataset, which presents a real-life fake news classification task broken down into two, three or six classes [nakamura2019r]. This dataset consists of more than one million posts on 22 different subreddits of the social networking platform Reddit and includes text and image modalities, supplemented by metadata about the posts and their authors. For the purposes of this study, a single binary data stream was prepared, in which consecutive texts were sorted accordingly to their creation timestamp. Of the entire dataset, $682,996$ multimodal samples were used (the rest have only one modality). The decision to limit to multimodal samples only is linked to the facilitation of extending the presented research to include the image modality. The data stream was divided into $2731$ data chunks, containing $250$ samples each. The resulting data stream is characterized by a dynamic imbalance ratio, which changes in successive batches to the point where, at about 3/4 of the stream length, a minority class transitions into a majority class. The changes in the prior probability of class membership are shown in Fig. 3.
@@
Experimental protocol & reproducibility All experiments were carried out using the Test-Then-Train protocol to guarantee a robust experimental evaluation, were implemented in Python and can be replicated using the publicly available GitHub repository [^1] https://github.com/w4k2/sentence-space-stream. Implementation of state-of-the-art algorithms were based on stream-learn [ksieniewicz2022stream], scikit-multiflow [montiel2018scikit], and PyTorch [paszke2017automatic] libraries. The classification quality evaluation of the algorithms was based on the standard metrics used in the task of imbalanced data classification, i.e. balanced accuracy score (bac), recall, specificity, precision, $F_{1}$ score, $Gmean$, and $Gmean_{s}$.
@@ 4.2 Experiment scenarios
Experiment 1 – Extraction methods The goal of Experiment 1 was to compare the performance of sss depending on the type of extractor used to obtain a representation for Sentence Space encoding. For this purpose, (aGloVe, (bMiniLM, (c) pretrained Word2Vec, and (dWord2Vec updated after each data chunk were compared with each other. The representation width depending on the extractor was $380$ for MiniLM and $300$ for GloVe and W2V. Based on the results, the feature extraction method used in subsequent experiments was selected.
Experiment 2 – Comparison with data stream classification algorithms In Experiment 2, the sss based on the extractor chosen in Experiment 1 was compared with state-of-the-art ensemble algorithms for imbalanced data stream classification. Among the methods mentioned in the literature review, (aHoeffding Tree with Hellinger split criterion (hf), (bLearn++.CDS (cds), (cLearn++ .NIE (nie), (dKappa Updated Ensemble (kue), and (eRobust Online Self-Adjusting Ensemble (rose) were selected as references. hf was used as the base classifier for all reference methods, and the maximum size of the classifier pool was set to $10$. The selection of methods, base classifier, and pool size was based on the literature [aguiar2023survey, cano2020kappa, cano2022rose]. The entire set of reference algorithms was compared with the sss depending on the approach used to extract features from the text. In addition to the extractors used in Experiment 1, tf-idf with unigrams and bigrams and $100$ features with top term frequency was employed here. A set of reference methods trained using the representation that provided the highest classification quality in terms of bac was selected for the last experiment.
@@ 4.3 Experiment 1 – Extraction methods
As can be seen in Figure 4, across all data chunks, the results of the first three methods – GloVe, MiniLM, and Word2Vec pretrained – are similar, with the only deviating method being Word2Vec partial-fit trained on each chunk. This discrepancy can be explained by the dictionary’s limitations resulting from training on successive data chunks containing only short texts. In contrast, MiniLM, Word2Vec pretrained (word2vec-google-news-300), and GloVe (glove-wiki-gigaword-300), were all built on larger volumes of data compared to the Word2Vec trained from scratch.
Following this experiment, it was decided to use the GloVe vectors for further research, which is a newer method than Word2Vec that processes additional global information and, at the same time – unlike MiniLM, dedicated to sequential processing – was designed to process single words. In addition, the MiniLM model requires more computational complexity when determining the representation vector for a given word, which for Word2Vec and Glove methods is limited only to reading values from the vector array. With equal effectiveness, the exclusion of MiniLM is fully justified. Thus, the results obtained in Experiment 1 made it possible to answer RQ1.
@@ 4.4 Experiment 2 – Comparison with data stream classification algorithms
The preliminary part of the main comparative experiment begins with analyzing the influence of feature space reduction, which was achieved using Principal Component Analysis (pca) – projecting the original embedding space down into $100$ features. Based on the achieved results, it can be concluded that the influence of this simple reduction on the results is mostly cosmetic, always laying in a one percent margin of a difference, so it is justified to reduce problem representation for canonical models since representation gets smaller while change in classification quality is negligible.=
The experimental evaluation results, taking into account the proposed sss approach, are presented in Figure 5, divided into four subfigures according to the extraction strategies used for canonical methods. Regardless of the extraction method, the nie algorithm performs the worst, maintaining bac at the level of a random classifier for almost the whole stream length. Only when using Glove and MiniLM embeddings does its balanced accuracy score rise slightly in the final phase of the stream, after the prior concept drift shown in Fig. 3.
@@
The second weakest method is Learning++CDS, which, for all extractors except MiniLM, clearly leans towards randomness. All analyzed reference solutions appear to be very sensitive to concept drift occurring around the chunk $2,000$. The ranking of kue, rose, and Hoeffding Tree algorithms depends on the extraction strategy used. However, their effectiveness’s overall distribution is similar, with a slight advantage for rose.
sss shows a noticeable advantage over all canonical solutions throughout the entire data stream, being the only one to maintain an average recognition efficiency of $80\%$ balanced accuracy score (thereby responding to RQ2). This observation is supported by the extended analysis of metrics in the form of a radar diagram (Figure 6), where an advantage of sss can be observed in each of the base and aggregated metrics used. The outlier specificity result for the nie method stems from its complete inability to learn in the analyzed problem environment, which for the dominant majority of processing time induces decisions towards only one of the problem classes, without a generalized relation to given bias.
@@
Among the analyzed extraction methods, MiniLM comes minimaly to the fore, as it was the only one that allowed the nie strategy to noticeably rise from the random classifier level in the final part of the stream, and the cds to compete with the rest of the algorithms. Accordingly, it was the MiniLM embeddings that were used for the canonical methods in Experiment 3.
@@ 4.5 Experiment 3 – Time complexity
The results of the third experiment – showing the time complexity for extracting, training, and testing the algorithm – are shown in Figure 7. As we can see, in the case of preprocessing, all reference methods show uniform computation time (lines overlap), and only the proposed approach deviates from this tendency and performs extraction faster. Additionally, it should be noted that this happens despite the time measurement considering the transition to image representation. The method owes it using the GloVe technique for sss, which proved to work best in Experiment 1. The reference methods, however, use the MiniLM transformer in accordance with the outcome of Experiment 2.
@@
More variability can be observed in the training and testing processes – in both, a single Hoeffding Tree processes the fastest. In turn, training in a single epoch is the slowest for rose, and prediction – for nie. It is also reflected in the accumulated time graph, which indicates that the only method ahead of sss regarding the entire processing time is the Hoeffding Tree. Still, it should be noted that a single classifier will always be faster than an ensemble.
Therefore, the proposed method is faster then the classifier ensembles used for data streams even though it is based on convolutional network and requires encoding from text into an image. At the same time, despite the lowest processing time compared to state-of-the-art ensemble algorithms, sss offers the highest generalization ability. All this means that sss can be successfully used in real-life batch-based data stream classification tasks, answering the RQ3.
5 Conclusions
The presented research work aimed to achieve two main goals. The first was to address the application of deep learning in the task of data stream classification, which is presented in the current literature as one of the main research areas in the need of further investigation. The second goal was to propose the use of sentence space, which is the equivalent of multi-dimensional encoding for text data, in a data stream classification task.
To realize the above goals, Streaming Sentence Space (sss) was proposed, which encodes the text found in individual data batches into discrete digital signals based on embeddings obtained through the GloVe technique. The resulting images are then classified using the ResNet-18 architecture, which, in order to reduce computational complexity, performs only a single training epoch on each data chunk. The developed approach was tested on the basis of computer experiments conducted on a real-life dynamically imbalanced data stream formed by chronologically ordering the texts contained in the fakeddit dataset. The results showed that sss, thanks to the inherent generalization ability of the convolutional neural network, is able to outperform the classification quality of state-of-the-art classifier ensemble methods dedicated for imbalanced data stream classification. In addition, sss exhibits lower time complexity than ensemble reference methods, which further encourages its use and contradicts the popular opinion that deep learning has too high time and computational complexity to be used for data stream analysis.
@@ 5 Conclusions
{credits}
5.0.1 Acknowledgements
This work was supported by the National Center for Research and Development within INFOSTRATEG program, number of application for funding: INFOSTRATEG-I/0019/2021-00.